OpenStudy (anonymous):

Input the equation of the given line in standard form. The line with m = -.5 and the point (-2, -3).

OpenStudy (azureilai):

standard form is y=mx+b To find the b, plug (-2,-3) into the point slope form (y-y1)=m(x-x1) so (y-(-3))=-.5(x-(-2)) y+3=-.5x-1 y=-.5x-4

jigglypuff314 (jigglypuff314):

first put what you know into the equation you were taught point-slope formula: y - y1 = m(x- x1) when given point (x1, y1) and m = slope or the slope-intercept formula: y = mx + b when m = slope then you solve it in such a way so that you get it into standard form: ax + by = c

OpenStudy (anonymous):

Well that answer I know.... I need the steps to the standard form equation....PLEASE :) :) :)

jigglypuff314 (jigglypuff314):

to get this to standard form (ax + by = c) you want to get the values with the x and y variables onto one side by themselves so y=-0.5x-4 add both sides by 0.5x +0.5x +0.5x 0.5x+y = 0 - 4 so your answer would be 0.5x + y = -4

OpenStudy (anonymous):

AWESOME THANKS>>> I JUST FINISHED IT AND I HAD THE SAME ANSWER>>>> thanks a million times.... :) :) :) :) :) :)

jigglypuff314 (jigglypuff314):

your welcome :)